Differentiate cube root of x plus cube root of x
1
Expert's answer
2019-11-18T12:22:21-0500

Differentiation


We need to differentiate the Function "\sqrt[3]{x + \sqrt[3]{x}}"


Solution:


We know,

"Derivative \space of \space x^n = n \times x^{n-1}"

Chain rule,

"Derivative of f(g(x)) = f ' (g(x)) \times g'(x) \times Derivative \space of \space (x)"


Let the given function as

"y = \sqrt[3]{x + \sqrt[3]{x}} = ( x + {x}^ {\frac {1}{3}})^{\frac {1}{3}}"

Now, Differentiate with respect to x,

"\frac {dy}{dx} = \frac {d}{dx} ( ( x + {x}^ {\frac {1}{3}})^{\frac {1}{3}})"


"= \frac {1}{3} \times ( x+ x^{\frac {1}{3}})^{(\frac {1}{3} - 1)} \times \frac {d}{dx} (x + x^{\frac{1}{3}})"

"= \frac {1}{3} \times ( x+ x^{\frac {1}{3}})^{(\frac {-2}{3} )} \times ( 1 + \frac {1}{3} \times x ^{\frac {-2}{3}})"




"= \frac {1}{9} \times ( x+ x^{\frac {1}{3}})^{(\frac {-2}{3} )} \times ( 3 + x ^{\frac {-2}{3}})"

Answer:

"\frac {dy}{dx} = = \frac {1}{9} \times ( x+ x^{\frac {1}{3}})^{(\frac {-2}{3} )} \times ( 3 + x ^{\frac {-2}{3}})"